If the price of a product decreases by 5% and the quantity demanded increases from 115 to 120, what is the price elasticity of demand for this product?
1
Expert's answer
2020-11-30T16:15:22-0500

Price Elasticity ="\frac{\Delta Q}{\Delta P}×100"


=115"-"120="-5%"


"\frac{-5}{115}×100" ="-4.3" =Quantity Demanded


Price Elasticity ="\frac{-4.3}{5}" ="-0.86"